# The Coding Limits of Robust Watermarking for Generative Models

### v1

@@ Abstract
We prove a sharp threshold for the robustness of cryptographic watermarking for generative models. This is achieved by introducing a coding abstraction, which we call messageless secret-key codes, that formalizes sufficient and necessary requirements of robust watermarking: soundness, tamper detection, and pseudorandomness. Thus, we establish that robustness has a precise limit: For binary outputs no scheme can survive if more than half of the encoded bits are modified, and for an alphabet of size $q$ the corresponding threshold is (11/q)(1-1/q) of the symbols.
Complementing this impossibility, we give explicit constructions that meet the bound up to a constant slack. For every $\delta>0$, assuming pseudorandom functions and access to a public counter, we build linear-time codes that tolerate up to $(1/2)(1-\delta)$ errors in the binary case and $(1-1/q)(1-\delta)$ errors in the $q$-ary case. Together with the lower bound, these yield the maximum robustness achievable under standard cryptographic assumptions.
@@
These results provide a complete characterization of robust watermarking, identifying the threshold at which robustness fails, constructions that achieve it, and an experimental confirmation that the threshold is already reached in practice.
@@ 1 Introduction
Watermarking is one of the most promising strategies for meeting this challenge [Exe23, Eur24]. The idea is to embed a secret, imperceptible pattern into model outputs at generation time, enabling later verification by anyone with the appropriate key. Ideally, a watermark should be robust, surviving even substantial adversarial editing, while also being undetectable to those without the key and leaving the original content unchanged in quality or meaning [Aar22, KGW+23, CG24].
@@
Recent work by Christ and Gunn [CG24] formalized cryptographic watermarking using pseudorandom error-correcting codes (PRCs), showing that such watermarks can be both undetectable and robust to significant amounts of tampering. But their results naturally lead to a deeper question: what are the true limits of robustness for any cryptographically grounded watermark? Is there a universal barrier that no watermarking scheme, however cleverly designed, can overcome?
@@
This question has been explored from two sides. On one side, the “Watermarks in the Sand” (WiTS) work [ZEF+23, ZEF+24] showed that any sufficiently robust watermark is, in principle, vulnerable: if an adversary has enough power, especially access to quality and perturbation oracles, it can eventually erase any watermark without degrading content quality. However, WiTS leaves open the concrete thresholds for robustness that efficient watermarking schemes can actually achieve. On the other side, PRC-based schemes demonstrate the power of current cryptographic techniques but do not establish whether their level of robustness is the best possible.
@@
In this paper, we bring these perspectives together. We introduce a conceptually simple abstraction, the messageless secret-key code, which crystallizes the core challenges and possibilities of cryptographic watermarking. Using this notion, we prove a tight information-theoretic threshold: no watermark, regardless of computational assumptions, can reliably survive if more than half of the encoded bits are altered in the binary case. More generally, for a $q$-ary alphabet, the limit is (11/q)(1-1/q) of the encoded symbols. Conversely, for any small constant $\delta$, we give explicit constructions that approach these limits, providing robustness up to just under half of the bits (for binary) or just under (11/q)(1-1/q) (for $q$-ary) of the symbols changed.
@@
To make our results concrete, we analyze a recent state-of-the-art PRC-based watermarking scheme for images by Gunn, Zhao, and Song [GZS25]. We show that a simple crop-and-resize operation, one that visually preserves the image but changes about half of its latent bits, is sufficient to reliably erase the watermark in practice. This demonstrates that the theoretical limit we establish is not just a worst-case artifact, but a real constraint in practical watermarking. This attack exposes a striking contrast between modalities: in text, watermark removal requires extensive edits that inevitably alter content, while for images, even a single benign transformation can erase the watermark without changing the image in any meaningful way. This makes the problem of robust watermarking in images uniquely delicate, as the watermark can disappear without a trace and without a visible cost.
@@
Our findings position the impossibility results of WiTS in a precise, quantitative framework: WiTS demonstrates that all robust watermarks are eventually breakable in the black-box oracle model, while our work identifies the exact numerical threshold at which robustness becomes fundamentally impossible for any PRC-style, cryptographically undetectable watermark.
@@
Looking forward, our results highlight that any significant increase in watermark robustness will require fundamentally new ideas, potentially leveraging semantic or structural features of content rather than cryptographic pseudorandomness alone.
@@ 1.1 Our Contributions
Our work provides a definitive and technically sharp characterization of the possibilities and limitations of cryptographic watermarking. Specifically, our contributions are as follows:
@@
    A Simplified Abstraction for Watermarking. We introduce and formalize messageless secret-key codes (also referred to as “zero-bit” in prior literature), a cryptographic primitive that captures the minimal requirements for watermarking generative models. These codes do not encode explicit messages but instead focus on pseudorandomness, soundness, and robust tamper-detection. We prove that messageless codes are not only sufficient but also necessary for cryptographic watermarking of generative AI outputs, even when the watermark detector has access to the prompt.
@@
    Optimal Constructions. We provide an explicit and efficient construction of messageless secret-key codes under the standard assumption of secure pseudorandom functions and the existence of an untamperable public counter. This construction achieves robustness up to nearly half the codeword symbols, as formalized below. Our construction also supports codes over larger alphabets.
    Theorem 1 (Optimal Messageless Code, Informal).
    For any constant $\delta>0$ and any alphabet of size $q\geq 2$, there exists a simple, explicit construction of a messageless secret-key code (using a PRF and untamperable counter) that achieves tamper-detection robustness against adversarial modification of fewer than the threshold $\alpha^{*}$ fraction of codeword symbols where
    $\alpha^{*}=\begin{cases}\frac{1}{2}(1-\delta),&\text{for binary alphabets;}\\ (1-1/q)(1-\delta),&\text{for $q$-ary alphabets.}\end{cases}$ (1)
@@
    Tight Information-Theoretic Limits. We prove a matching lower bound: no messageless secret-key code (and hence no cryptographic watermarking scheme, even with negligible soundness and tamper-detection error) can robustly detect tampering if an adversary can modify more than the threshold $\alpha^{*}$ (as defined in (1) above) fraction of the symbols, for any constant $\delta>0$. This impossibility holds even if the watermark detector is provided the prompt, and applies to both secret-key and public-key versions of the primitive.
@@
    Optimality of PRC-Based Watermarking. Our upper and lower bounds show that the robustness achieved by the PRC watermarking schemes of Christ and Gunn [CG24] is in fact information-theoretically optimal. Any cryptographic watermark (binary or $q$-ary, secret-key or public-key) cannot exceed these limits.
  5. 5.
    Concrete Attack. We demonstrate that these theoretical limits are tight and immediately relevant in practice. Specifically, we show that a simple crop-and-resize transformation modifies enough underlying latent bits to consistently and silently remove the state-of-the-art PRC watermark [GZS25], without visible quality loss.
    Earlier work had already shown that watermarks can in principle be erased. WiTS established this with oracle access, showing that by repeatedly regenerating high-quality outputs an adversary can eventually reach an unmarked image [ZEF+23, ZEF+24]. More recently, the UnMarker attack [KH24] demonstrated removal in a fully black-box setting, but only by perturbing essentially every pixel through computationally heavy spectral optimization, leaving behind small but measurable degradation.
    What distinguishes our result is that no such machinery is required. A single, routine crop-and-resize suffices to drive the system exactly to the threshold where our lower bound guarantees failure. In this sense, the limit is not just a theoretical boundary but a practical one, realized by the most ordinary of edits.
Collectively, these results establish the precise information-theoretic boundary for cryptographic watermark robustness, even under generous adversarial models. They show both what is achievable, and what is fundamentally impossible, with cryptographic watermarking alone.
@@ 1.2 Related work
Watermarking refers to the process of embedding a signal in generated content such as images [BOD95, CMB+07, HD17, RP97, ZKJFF18], text [ARC+01, ARH+03], audio [Arn00, BTH96], or video that is imperceptible to humans but algorithmically detectable. Its primary use is to enable attribution or provenance without significantly degrading content quality. The study of watermarking spans decades and includes both practical systems and formal analyses, with results on constructing watermarking schemes as well as breaking or proving impossibility. With the rise of modern generative models, watermarking has become a sharper test of the trade-offs between imperceptibility, robustness, and efficiency.
@@
Watermarking methods are commonly classified based on when the watermark is applied: either after generation (post-processing) or during the generation process itself (in-processing). Post-processing schemes embed the watermark into the output after it has been generated. These approaches often degrade output quality [CMB+07, WWZ+22, ADR+24]. In-processing schemes, by contrast, embed the watermark during content generation. They were first introduced in image generators [YSAF21], and later adapted to large language models (LLMs) [KGW+23]. Image watermarking techniques include modifying weights through pretraining or fine-tuning [FXTB22, FCJ+23, ZZXP23, ZPD+23], or intervening in the sampling trajectory of diffusion models [WKGG23]. In the LLM setting, most methods operate at inference time by biasing token selection in a way that encodes a hidden message [KGW+23, CGZ24]. Some schemes further ensure that the overall distribution over outputs remains statistically indistinguishable from unmarked outputs [KTHL24], preserving quality.
A more recent line of work leverages cryptographic and coding-theoretic tools to design undetectable watermarking schemes. One direction [CG24] proceeds via pseudorandom error-correcting codes (PRCs) as the watermark signals. These codes can be constructed using low-density parity-check (LDPC) codes [CG24, AAC+25, GG24] and allow for robust detection under edits. PRC-based watermarking was recently applied to images [GZS25]. The main limitation [GZS25] is that although the underlying PRC is robust to modifications on a bounded number of bits, the resulting PRC-based watermarking is robust only to independent changes, limiting its applicability in real-world scenarios. In parallel, Fairoze et al. [FGJ+25] propose a public watermarking scheme that uses digital signatures, allowing third parties to verify the watermark and ensuring robustness as long as a small window of the generated content remains unmodified.
@@
Alongside construction efforts, recent work has explored the limitations of existing watermarking schemes. In text, the most prominent attack is paraphrasing [CGLS25, KSK+23, ZEF+23, ZEF+24], rewriting the watermarked content while preserving its meaning, using language models, translation tools, or manual edits. These attacks often succeed partially, but at the cost of degraded output quality. In the vision domain, a range of attacks aim to remove watermarks by adding noise to the image or latent representation, or by using optimization techniques to remove the watermark [LDFK24, SSR+24, ZZS+24]. Many of these attacks have limitations. For example, Zhao et al. [ZZS+24] focus on post-hoc watermarking methods, while Saberi et al. [SSR+24] require either white-box access to the detector or access to many watermarked and unwatermarked examples.
A more general impossibility result by Zhang et al. [ZEF+24, ZEF+23] shows that constructing robust watermarking schemes is impossible under certain conditions. They model generated outputs as nodes in a graph and prove that if an adversary has access to two oracles (a quality oracle that tells the adversary whether the output is still valid, and a perturbation oracle that allows small edits preserving semantics) then the adversary can walk through the graph until the watermark is erased without degrading content quality. This result does not specify an exact condition for when removal occurs, only that eventually such a walk succeeds.
In contrast, our result makes no oracle assumptions and provides a concrete lower bound. We show that if an adversary is willing to change more than half of the bits, then the watermark is surely removed. Prior work demonstrates that robust watermarking cannot be sustained indefinitely. Our result identifies the precise threshold at which robustness fails.
@@ 2.1 Notation
Small letters (such as $x$) denote individual objects or values; calligraphic letters (such as $\mathcal{X}$) denote sets; sans-serif letters (such as $\mathsf{A}$) denote algorithms. For a string $x\in\{0,1\}^{*}$, $|x|$ denotes its length. For a set $\mathcal{X}$, $|\mathcal{X}|$ is its cardinality. If $x$ is chosen uniformly at random from $\mathcal{X}$, we write $x{\>{\leftarrow{\hskip-3.0pt\raisebox{0.75pt}{$\scriptscriptstyle\$$}}}\>}\mathcal{X}$. The Hamming distance $\mathrm{dist}(x,x^{\prime})$ between two strings $x$ and $x^{\prime}$ (over an alphabet $\Sigma$}) is the number of positions where they differ.
For a deterministic algorithm $\mathsf{A}$, $y=\mathsf{A}(x)$ means that $y$ is the output of $\mathsf{A}$ on input $x$. For a randomized algorithm $\mathsf{A}$, we write $y{\>{\leftarrow{\hskip-3.0pt\raisebox{0.75pt}{$\scriptscriptstyle\$$}}}\>}\mathsf{A}(x)$ to indicate $y$ is sampled from the output distribution of $\mathsf{A}$ on input $x$. Alternatively, $y=\mathsf{A}(x;r)$ denotes the output when $\mathsf{A}$ runs on $x$ with randomness $r$. An algorithm $\mathsf{A}$ is called probabilistic polynomial-time (PPT) if it is randomized and always halts in time polynomial in $|x|$.

### v2

@@ Abstract
We ask a basic question about cryptographic watermarking for generative models: to what extent can a watermark remain reliable when an adversary is allowed to corrupt the encoded signal? To study this question, we introduce a minimal coding abstraction that we call a zero-bit tamper-detection code. This is a secret-key procedure that samples a pseudorandom codeword and, given a candidate word, decides whether it should be treated as unmarked content or as the result of tampering with a valid codeword. It captures the two core requirements of robust watermarking: soundness and tamper detection.
Within this abstraction we prove a sharp unconditional limit on robustness to independent symbol corruption. For an alphabet of size $q$, there is a critical corruption rate of 11/q1-1/q such that no scheme with soundness, even relaxed to allow a fixed constant false positive probability on random content, can reliably detect tampering once an adversary can change more than this fraction of symbols. In particular, in the binary case no cryptographic watermark can remain robust if more than half of the encoded bits are modified. We also show that this threshold is tight by giving simple information-theoretic constructions that achieve soundness and tamper detection for all strictly smaller corruption rates.
@@ 1 Introduction
Watermarking is one of the most promising strategies for meeting this challenge [US23, EU24]. The idea is to embed a secret, imperceptible pattern into model outputs at generation time, enabling later verification by anyone with the appropriate key. Ideally, a watermark should be robust, surviving even substantial adversarial editing, while also being undetectable to those without the key and leaving the original content unchanged in quality or meaning [Aaronson22, Kirchenbauer_23, C:ChrGun24].
@@
Recent work by Christ and Gunn [C:ChrGun24] formalized cryptographic watermarking using pseudorandom error-correcting codes (PRCs), showing that such watermarks can be both undetectable and robust to significant amounts of tampering. But their results naturally lead to a deeper question: within such cryptographically grounded, coding-based frameworks, what are the true limits of robustness? Is there an unconditional barrier that no watermarking scheme of this form can overcome?
@@
This question has been explored from two sides. On one side, the “Watermarks in the Sand” (WiTS) work [zhang2023watermarks, zhang2024watermarks] showed that any sufficiently robust watermark is, in principle, vulnerable: if an adversary has enough power, especially access to quality and perturbation oracles, it can eventually erase any watermark without degrading content quality. However, WiTS leaves open the concrete thresholds for robustness that efficient watermarking schemes can actually achieve. On the other side, PRC-based schemes demonstrate the power of current cryptographic techniques but do not establish whether their level of robustness is the best possible in the coding sense.
@@
In this paper, we bring these perspectives together. We introduce a conceptually simple abstraction, the zero-bit tamper-detection code, which we use as a minimal model for cryptographic watermarking based on pseudorandom codes. This notion can be viewed as the zero-bit specialization of a PRC equipped with an explicit tamper-detection predicate, and it crystallizes the core challenges and possibilities of such watermarking schemes. Using this notion, we prove a tight unconditional threshold on robustness to independent symbol corruption in this model: no scheme realizing a zero-bit tamper-detection code can reliably survive if more than half of the encoded bits are altered in the binary case. More generally, for a $q$-ary alphabet the corresponding limit is (11/q)(1-1/q) of the encoded symbols. Conversely, for any small constant $\delta$, we give explicit constructions that approach these limits, providing robustness up to just under half of the bits (for binary) or just under (11/q)(1-1/q) (for $q$-ary) of the symbols changed.
@@
To make our results concrete, we analyze a recent state-of-the-art PRC-based watermarking scheme for images by Gunn, Zhao, and Song [gunn2025undetectable]. We show that a simple crop-and-resize operation, one that visually preserves the image but changes about half of its latent bits, is sufficient to consistently prevent belief-propagation decoding from recovering the codeword. This demonstrates that the theoretical limit we establish is not just a worst-case artifact of the model, but a real constraint in at least one practical watermarking pipeline.
This attack exposes a striking contrast between modalities: in text, watermark removal typically requires extensive edits that inevitably alter content, while for images, even a single benign transformation can erase the watermark without changing the image in any meaningful way. This makes the problem of robust watermarking in images uniquely delicate, as the watermark can disappear without a trace and without a visible cost.
@@
Our findings position the impossibility results of WiTS in a precise, quantitative framework for PRC-style cryptographic watermarks: WiTS demonstrates that all robust watermarks are eventually breakable in a powerful black-box oracle model, while our work identifies a concrete numerical threshold at which robustness becomes unconditionally impossible for schemes that can be modeled as zero-bit tamper-detection codes under independent symbol corruption.
@@
Looking forward, our results suggest that any significant increase in robustness within this PRC-based, cryptographic framework will require fundamentally new ideas, potentially leveraging semantic or structural features of content rather than cryptographic pseudorandomness alone.
@@ 1.1 Our Contributions
Our work provides a clean and technically precise view of what cryptographic watermarking based on pseudorandom codes can and cannot achieve. Specifically, our contributions are as follows:
@@
    A simplified abstraction for PRC-style watermarking. We introduce and formalize zero-bit tamper-detection codes, a secret-key primitive that we use as a minimal model for cryptographic watermarking based on pseudorandom codes. These codes do not encode explicit messages but instead focus on two properties: soundness and tamper detection. They can be viewed as the zero-bit specialization of pseudorandom error-correcting codes equipped with an explicit tamper-detection predicate (and without psedorandom codewords). We show how any watermarking scheme that is sound and robust for a given family of tampering functions induces a zero-bit tamper-detection code with the same guarantees, even when the watermark detector depends on the prompt. In this sense, zero-bit tamper-detection codes capture the core coding-theoretic requirements of PRC-based watermarking for generative models.
@@
    Tight unconditional limits for watermarking schemes. Within this abstraction we establish a sharp limit on robustness to independent symbol tampering. For an alphabet of size $q$ there is a critical corruption rate
    $\alpha^{*}=1-\frac{1}{q}$
    such that no zero-bit tamper-detection code can simultaneously satisfy soundness and tamper detection for independent tampering at rates exceeding $\alpha^{*}$ by any fixed constant slack. In particular, in the binary case ($q=2$) no such code can remain both sound and robust once more than half of the encoded bits are modified. Conversely, we give a simple information-theoretic construction that achieves soundness and tamper detection for all corruption rates strictly below $\alpha^{*}$, up to an arbitrarily small constant slack.
    We further show that this barrier persists even if soundness is weakened to constant-bounded soundness. For any constant $c\in(0,1]$, once an adversary can corrupt more than an $(\alpha^{*}+\varepsilon)$ fraction of symbols (for any fixed $\varepsilon>0$), any code that is sound on all but a $c$ fraction of random strings must incur tamper-detection error close to 1c1-c at that corruption level. In particular, even very weak schemes that behave like a coin flip on random inputs (for example, $c=1/2$) cannot robustly detect tampering beyond this threshold.
@@
    Implications for PRC-based watermarking. By relating zero-bit tamper-detection codes to pseudorandom error-correcting codes, and using the reduction from watermarking schemes to tamper-detection codes, we obtain direct consequences for PRC-based watermarking. In particular, for the independent symbol tampering model considered by Christ and Gunn [C:ChrGun24], our lower bound shows that their robustness guarantees are optimal: no cryptographic watermark (binary or $q$-ary, secret-key or public-key) whose robustness can be expressed through such a coding layer can tolerate a larger fraction of symbol errors in this model.
@@
    A concrete attack illustrating the limit. Finally, we show that the coding-theoretic limit is numerically relevant for at least one state-of-the-art watermarking pipeline. We analyze the PRC-based image watermark for Stable Diffusion proposed by Gunn, Zhao, and Song [gunn2025undetectable], and we study the effect of a simple crop-and-resize transformation. This transformation preserves the visual appearance of the image yet induces latent sign errors concentrated near fifty percent and consistently causes belief-propagation decoding to fail, thereby erasing the watermark.
    Earlier work had already established that watermarks can, in principle, be removed. WiTS shows this using oracle access and iterative regeneration [zhang2023watermarks, zhang2024watermarks], while UnMarker demonstrates removal in a fully black-box setting using optimization-based spectral attacks that slightly degrade the image [unmarker2025]. Our case study differs in that it reaches essentially the same corruption regime using only a single, routine image operation, with no oracle access and no optimization. In this sense, it provides a concrete example where the unconditional robustness threshold arises from an ordinary edit rather than a highly tailored attack.
@@ 1.2 Related work
Watermarking refers to the process of embedding a signal in generated content such as images [boland_95, cox_et_al_07, Hayes_et_al_17, Ruanaidh_1997, Zhu_18], text [text_1, text_2], audio [arnold_00, Boney_96], or video that is imperceptible to humans but algorithmically detectable. Its primary use is to enable attribution or provenance without significantly degrading content quality. The study of watermarking spans decades and includes both practical systems and formal analyses, with results on constructing watermarking schemes as well as breaking or proving impossibility. With the rise of modern generative models, watermarking has become a sharper test of the trade-offs between imperceptibility, robustness, and efficiency.
@@
Watermarking methods are commonly classified based on when the watermark is applied: either after generation (post-processing) or during the generation process itself (in-processing). Post-processing schemes embed the watermark into the output after it has been generated. These approaches often degrade output quality [cox_et_al_07, Wan_22, An_24]. In-processing schemes, by contrast, embed the watermark during content generation. They were first introduced in image generators [Yu_21], and later adapted to large language models (LLMs) [Kirchenbauer_23]. Image watermarking techniques include modifying weights through pretraining or fine-tuning [Fei_22, Fernandez_23, Zeng_23, Zhao_23c], or intervening in the sampling trajectory of diffusion models [wen_2023]. In the LLM setting, most methods operate at inference time by biasing token selection in a way that encodes a hidden message [Kirchenbauer_23, Christ_23]. Some schemes further ensure that the overall distribution over outputs remains statistically indistinguishable from unmarked outputs [kuditipudi_24], preserving quality.
A more recent line of work leverages cryptographic and coding-theoretic tools to design undetectable watermarking schemes. One direction [C:ChrGun24] proceeds via pseudorandom error-correcting codes (PRCs) as the watermark signals. These codes can be constructed using low-density parity-check (LDPC) codes [C:ChrGun24, STOC:AACDG25, ghentiyala_24] and allow for robust detection under edits. PRC-based watermarking was recently applied to images [gunn2025undetectable]. The main limitation [gunn2025undetectable] is that although the underlying PRC is robust to modifications on a bounded number of bits, the resulting PRC-based watermarking is robust only to independent changes, limiting its applicability in real-world scenarios. In parallel, a recent public watermarking scheme uses digital signatures, allowing third parties to verify the watermark and ensuring robustness as long as a small window of the generated content remains unmodified [garg23].
@@
Alongside construction efforts, recent work has explored the limitations of existing watermarking schemes. In text, the most prominent attack is paraphrasing [paraphrasing_attack_1, paraphrasing_attack_2, zhang2023watermarks, zhang2024watermarks], rewriting the watermarked content while preserving its meaning, using language models, translation tools, or manual edits. These attacks often succeed partially, but at the cost of degraded output quality. In the vision domain, a range of attacks aim to remove watermarks by adding noise to the image or latent representation, or by using optimization techniques to remove the watermark [lukas_24, saberi_24, Zhao_24b]. Many of these attacks have limitations. For example, Zhao et al. [Zhao_24b] focus on post-hoc watermarking methods, while Saberi et al. [saberi_24] require either white-box access to the detector or access to many watermarked and unwatermarked examples.
A more general impossibility result by Zhang et al. [zhang2024watermarks, zhang2023watermarks] shows that constructing robust watermarking schemes is impossible under certain conditions. They model generated outputs as nodes in a graph and prove that if an adversary has access to two oracles (a quality oracle that tells the adversary whether the output is still valid, and a perturbation oracle that allows small edits preserving semantics) then the adversary can walk through the graph until the watermark is erased without degrading content quality. This result does not specify an exact condition for when removal occurs, only that eventually such a walk succeeds.
In contrast, our result makes no oracle assumptions and provides a concrete lower bound in an independent symbol corruption model. We show that if more than half of the encoded bits are effectively randomized, then no watermark can remain both sound and robust. Prior work demonstrates that robust watermarking cannot be sustained indefinitely. Our result identifies a precise corruption threshold in this setting beyond which robustness is unconditionally impossible.
@@ 2.1 Notation
Small letters (such as $x$) denote individual objects or values; calligraphic letters (such as $\mathcal{X}$) denote sets; sans-serif letters (such as $\mathsf{A}$) denote algorithms. For a string $x\in\{0,1\}^{*}$, $|x|$ denotes its length. For a set $\mathcal{X}$, $|\mathcal{X}|$ is its cardinality. If $x$ is chosen uniformly at random from $\mathcal{X}$, we write $x{\>{\leftarrow{\hskip-3.0pt\raisebox{0.75pt}{$\scriptscriptstyle\mathdollar$}}}\>}\mathcal{X}$. The Hamming distance $\mathrm{dist}(x,x^{\prime})$ between two strings $x$ and $x^{\prime}$ (over an alphabet $\Sigma$}) is the number of positions where they differ.
For a deterministic algorithm $\mathsf{A}$, $y=\mathsf{A}(x)$ means that $y$ is the output of $\mathsf{A}$ on input $x$. For a randomized algorithm $\mathsf{A}$, we write $y{\>{\leftarrow{\hskip-3.0pt\raisebox{0.75pt}{$\scriptscriptstyle\mathdollar$}}}\>}\mathsf{A}(x)$ to indicate $y$ is sampled from the output distribution of $\mathsf{A}$ on input $x$. Alternatively, $y=\mathsf{A}(x;r)$ denotes the output when $\mathsf{A}$ runs on $x$ with randomness $r$. An algorithm $\mathsf{A}$ is called probabilistic polynomial-time (PPT) if it is randomized and always halts in time polynomial in $|x|$.